#7 (50 pts) Use an appropriate method to solve the given system 1 8 X= +
Answer #1

T9 102? het Ä 8 A = - - ľ -O -- ht to the Then AK the system be x- Axt fett, where fit= let Characteristic ear of a in 8 It3t Fundamental matris in 4 e p(t) -ze-3t e 34 e-3t e-3t adi [olts] =( e 3t -3t -ti ze he 13t deptolt 4e3t 2e-3t 4 e3t det [plSo, general Jorn in X=Xn (A typity =( - Zet_btet 6 24 et 3 het Or, x= Ge (4)+422=3+ (?) il - 3 e tbt et + -3 3 et 4 et - 24 (